OpenStudy (anonymous):

5x/7-1/14=9/28 solve for x

OpenStudy (anonymous):

well I did this and I came up with x=11/20

OpenStudy (anonymous):

to arrive at x= 11/20 you would start from the left hand side of the equation meaning the 5x/7-1/14 would be the first thing you want to work with it's basically to decide on if it can be simplified you'll multiply 5 and x then multiply x and 1. So it'll be 5*x which comes out to be 5x then you divide x by 1 which will be x. you multiply 5*x/7 and you get 5/7x so basically it's 5x over 7 (sorry if I am confusing you I am not trying to) 1/14 is 1 over 14 5*x/7-1/14 is 5x/7-1/14 so basically the left hand side is 5x/7-1/14 and thus began on the right hand side of the equation 9/28 it'll just be 9/28 you don't have to really do anything to it and then we set up our equation because the left hand side was simplified the equation turns out to be 5x/7-1/14=9/28 you can move -1/14 to the right hand side by adding 1/14 to both sides it'll be -1/14+1/14 which equals zero making the answer 5x/7 to the right hand side it'll be 9/28+1/14=11/28 and the equation will be 5x/7=11/28 in order to isolate the x you have to divide both sides of the equation by the other variables. so you have to divide both sides by 5/7 first divide x by 1 which is x 5x/7 divided by 5/7 =x 11/28 divided by 5/7 = 11/20 so x is 11/20 hope that helped sorry if I was confusing you can ask me more if you want.